How do you find x+y=5
x-y=3

Answer by Alan3354(69443)

How do you find x+y=5
x-y=3
---------
Find what?
==========
Solve for x & y
x+y=5
x-y=3
------- Add
2x = 8
x = 4
------------
x+y=5
4+y = 5
y = 1